40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 20 11:57:17 2015, ericq, Update, General, Arm Locking recovered 
    Reply  Wed Jul 22 15:09:40 2015, ericq, Update, General, IMC work 
       Reply  Mon Aug 10 11:57:45 2015, ericq, Update, General, IMC work 
Message ID: 11430     Entry time: Mon Jul 20 11:57:17 2015     Reply to this: 11436
Author: ericq 
Type: Update 
Category: General 
Subject: Arm Locking recovered 

The interferometer is warming up!

I had some issues locking the IMC at first. It turned out that the MC3 side OSEM signal wasn't getting to the ADC. A satellite box sqush fixed it. 

I touched up the PMC alignment; the best I could do is 0.75V, probably due to the AOM being in place. 

I haven't touched the WFS offsets, but the current ones seem to be doing ok. I'll touch them up tonight when the seismic activity has calmed. 

I made some changes to the state of the PZT/PC crossover gain in the mcdown script, resulting in the IMC catching lock quicker. 

Thankfully, the tip tilt pointing stayed good during the upgrade. I barely had to touch the ETM alignment to lock the arms. ETMX is showing some errant motion, though... 

ELOG V3.1.3-